OpenStudy (tia_sweetheart17):

graph the linear equation. y=-2/5x-3

OpenStudy (anonymous):

to graph a linear equation, all you need are 2 points on the line that you can get from this equation, to get two points on the line just choose any 2 values for y and solve for x when y are the values you chose

OpenStudy (anonymous):

also, since we know that the y-intercept is -3, then a point right there is (0, -3) so all we need is one more point :)

OpenStudy (anonymous):

because y = mx + c and if you compare this to y = -2/5x -3. you can tell c = -3. and c is the y - intercept

OpenStudy (anonymous):

so find another point on the line and ill graph it for you :)

OpenStudy (tia_sweetheart17):

is 2 alright?

OpenStudy (anonymous):

if 2 is a value you want to choose then sure, any value is gud, my substitute x for 2 and solve for y :)

OpenStudy (tia_sweetheart17):

is it -4/5?

OpenStudy (anonymous):

sry i didnt get a notification. gud so far so y = -2/5 (2) -3 y = -4/5 - 3 so subtract -3 from -4/5 :)

OpenStudy (tia_sweetheart17):

11/5